OpenStudy (anonymous):

ATT James: An agricultural worker in Uganda is planting clover to increase the number of bees making their home in the region. There are a 100 bees in the region naturally, and for every acre put under clover 20 more bees are found in the region

OpenStudy (anonymous):

Draw a graph of the total # of bees as a function of x, the number of acres devoted to the clover

OpenStudy (anonymous):

What type of function is this?

OpenStudy (agreene):

y=20x+100 where y is bees and x is acres of clover.

OpenStudy (anonymous):

I thought it was a line but

OpenStudy (turingtest):

total bees intially=100 bee incresease per acre=20 acres=x total bees=y y=100+20x I was almost done anyway, so I posted

OpenStudy (anonymous):

K the question continues

OpenStudy (anonymous):

Explain both geometrically and algaebraically the shape of the graoh of

OpenStudy (anonymous):

The marginal rate of increase in the number of bees with acres of clover

OpenStudy (anonymous):

which is N'(x)

OpenStudy (turingtest):

You pretty much the geometric part did aleady, but you may as well talk about the intercepts for the algebraic interpretation.

OpenStudy (anonymous):

well the marginal rate wld be a straight line like y= whtvr

OpenStudy (turingtest):

if N(x)=20x+100 N'(x)=20 that is the marginal rate

OpenStudy (anonymous):

oh ok

OpenStudy (turingtest):

so it is a straight line, but a horizontal one obviously.

OpenStudy (anonymous):

and then you gotta find the shape of the graph geometrically and algebraically of the number of bees per acre which is N(x)/x

OpenStudy (turingtest):

well write it out, what is N(x)/x ?

OpenStudy (anonymous):

(100+20x)/x

OpenStudy (turingtest):

simplify...

OpenStudy (anonymous):

20(5+x)/x

OpenStudy (turingtest):

I meant divide to get rid of an x N(x)/x=20+100/x

OpenStudy (turingtest):

so what is the identifying part of this equation in terms of the shape of the graph?

OpenStudy (anonymous):

100/x

OpenStudy (turingtest):

good, and I assume you know what this looks like, yes? You can also get the y-intercepts from your eqaution as well. What do you get for that?

OpenStudy (turingtest):

sorry, x-intercepts

OpenStudy (anonymous):

-5

OpenStudy (turingtest):

exactly :) I don't know much to say about the shape besides that. here's a pic http://www.wolframalpha.com/input/?i=20%2B100%2Fx

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have a question

OpenStudy (turingtest):

shoot

OpenStudy (anonymous):

what iss the difference btwn the derivative and the average number of bees per acre

OpenStudy (anonymous):

one is the average rate of change and the other is the instantaneous rate of change?

OpenStudy (turingtest):

No, they reflect different things. Bees per acre, which is found by simply dividing by x, gives the density of bees in the resulting fields of clover planted. The derivative tell how many MORE bees you get for each acre planted. The derivative reflects a rate of change, the number of bees per acre is a density function.

OpenStudy (anonymous):

thanks that clarified everything

OpenStudy (turingtest):

anytime :)